Price Corp. is considering selling to a group of new customers and creating new annual sales of $180,000. 4% will be uncollectible. The collection cost on these accounts is 2% of new sales, the cost of producing and selling is 76% of sales and the firm is in the 29% tax bracket. What is the profit on new sales?
Answer #1
Sales $ 180,000
Less:
Cost of producing and selling3 $ 136,800 =180000*76%
Bad debts $      7,200 =180000*4%
Collection cost $      3,600 =180000*2%
Profit before tax $    32,400
Less: Tax $      9,396 =32400*29%
Profit on sales $    23,004